# Computational Obfuscations and Random Oracles for Derandomizing Asynchronous Consensus

### v1

@@ Abstract
A method for converting an asynchronous randomized consensus to a deterministic asynchronous consensus is presented. The method uses program computation obfuscation (e.g., [5, 12]) and a random oracle (e.g., [9, 6]), assuming a computationally bounded scheduler, to overcome the impossibility result of Fischer, Lynch, and Paterson [10].
Two stages are combined, in the first, a new form of computational program obfuscation implemented by post-quantum cryptographic hash functions (see e.g., [12]) is introduced, employing time lock puzzles (see e.g., [19, 1]) to imply a computational gap between the consensus participants and the (imaginary adversarial) scheduler. In the second stage, a random oracle is implemented by using a post-quantum cryptographic hash function that allows each process to harvest pseudo-randomization from the (cleartext) program and a (consensus) round (variable) and, in turn, implies the completion of the consensus in the presence of a computationally bounded scheduler.
@@ 1 Introduction
Asynchronous consensus is a building block in structuring replicated state machines and efficient proof of authority/stake Blockchains, e.g., [8].
@@
Unfortunately, the celebrated result of Fischer, Lynch, and Paterson [10] proves that there is no deterministic algorithm that can yield an asynchronous consensus. Paxos [14] succeeded in ensuring the consistency (all participants decide on the same value) and validity (the decided value appears in at least one of the inputs) of the distributed decision but sacrifices the liveness as an adversarial scheduler can keep the system undecided forever.
Randomization comes to the rescue (e.g., [7]). We investigate the possibility of using a random oracle [9] to derandomize a randomized asynchronous consensus using a post-quantum one-way function.
Structuring the random oracle led us to use post-quantum time lock puzzles (see, e.g., [19, 1] and the references therein) for efficiently structuring obfuscated programs (see, e.g., [12] and the references therein). Intuitively, to obfuscate the program, we computationally hide (by means of time lock puzzles) an important binary truth table of a threshold function, where the adversary scheduler needs to recover the threshold index, while the participants need only recover the value in two indices. The result is deterministic asynchronous consensus in the presence of a computationally bounded adversary.
Overview.
The system settings are described in the next section. The distributed consensus task is described in Section 3, including the impossibility result to achieve consistency on the decided (same) output (agreement), output that reflects at least one of the inputs (validity) and is chosen to be output in a finite number of steps (termination). Then, randomization is introduced, assisting in solving the challenging consensus problem. Section 5 describes a fundamental building block in randomized consensus, namely the conciliator. A conciliator is a Monte Carlo consensus object, based on approximate agreement, that guarantees validity and termination in all executions but guarantees agreement only with some probability. An implementation of a conciliator is described in Section 5. Obfuscation is used to yield a (tunable) computation logarithmic factor between the computation the scheduler has to invest (with relation to the participants) to avoid agreement in the conciliator output. An additional randomized consensus technique is used to overcome the unfortunate case where the conciliator outputs disagree. In this stage, the use of a random oracle forces the adversarial scheduler to invest more computation than each participant, alas risking a successful consensus. Thus, intuitively, the computation budget of the scheduler is exhausted in order to reach a consensus.
Implementations of the obfuscation and random oracle appear in Section 6 and Section 8, respectively.
2 The system
Processes
@@ 2 The system
We define $p_{1},...,p_{n}$ to be a set of asynchronous processes. These processes communicate by reading and writing to shared objects.
Configuration
We define a configuration to be the state of the local memory of all $n$ processes $p_{1},\ldots,p_{n}$, including their program counters, together with the content of all shared objects.
Scheduler
We define a scheduler as an adversary that can decide which of the processes to activate next. This adversary can also follow the process program as dictated by the chosen process program counter and perform local computations that end with either a read operation from a shared object or a write operation to a shared object. The scheduler has complete information, namely, the scheduler knows the content of the shared objects, the local memory of each process, and the algorithm they execute.
Random Oracle
@@
We define a random oracle $H$ to be an idealized version of the cryptographic hash function, that is, a black box that can be queried with input $x$ to return output $y$. The box is consistent; that is, any input $x$ will always output the same $y$. The box is uniformly random, which means the probability of getting output $y$ for an input $x$ that was not already queried is uniform. Everyone has access to this black box and can query it, but its internal workings are unknown (See Chapter 5 of [13]).
3 The consensus task
@@
4 Bound on the cost of finding a bad execution
@@
As defined by Fischer, Lynch, and Paterson [10], a configuration $C$ is bivalent if there is a path from it to a configuration where a process decides on the value $1$, and there is also a path from it to another configuration where a process decides on the value 00. Similarly, a configuration $C$ is univalent if all the paths from $C$ lead to configurations with an identical decision value. Because of the agreement property, any configuration in which some process has decided is necessarily univalent.
@@
The FLP proof, in both its original form [10], and its extension to shared memory by Loui and Abu-Amara [15], gives an explicit procedure for constructing an infinite non-terminating execution for any protocol that satisfies agreement and validity, by starting in a bivalent configuration and always choosing bivalent successor configurations. However, the computational cost of this construction is not considered. We would like to find an adversary strategy that allows it to prevent a protocol from terminating without having to explore the (possibly infinite!) space of all executions.
@@
The time complexity of computing a non-terminating, and thus infinite, execution is necessarily infinite. So, we will look at the more tractable problem of computing an extension for some fixed number of steps $s$. The goal is to replace the abstract adversary function with an adversary program that takes as input the starting configuration of the system and the programs used by the processes to choose their next operations, and, in the case of agreement, their outputs.
It is convenient to model this problem in terms of a weakened version of consensus where termination within a fixed number of steps and validity are always required, but agreement may fail. This is similar to the conciliator abstraction of [3], with the difference that we can no longer assign a probability to failure but must instead rely on failing executions being computationally difficult to find.
@@
Formally, let us define the following problem. We will specify the behavior of each process using a random-access machine program $P$ that implements a function $Q×V\textrightarrow Q×A$, where $Q$ is the set of states of the process, $V$ is the set of possible return values of operations, and $A$ is the next operation to execute. An adversary is a program in the same model that takes as input the step bound $s$, a sequence of programs $\left\langle{P_{1},\dots,P_{n}}\right\rangle$ and initial states $q_{1},\dots,q_{n}$, and outputs a sequence of process ids specifying the order in which the processes take steps. We assume that each program $P_{i}$ has two special return actions $r_{0}$ and $r_{1}$, that it must choose between after exactly $s$ steps; these provide the output of each process in the consensus protocol. The adversary wins if any of these outputs disagree.
@@
Given inputs $P_{1},\dots,P_{n}$ representing processes in a shared-memory system where, in all executions, (a) each process $p_{i}$ outputs a decision value within $s$ steps, and (b) each process $p_{i}$ uses a total of at most $t$ time to compute its transitions, there is an adversary program that computes within $O(st)$ time a schedule that causes some pair of processes to output different decision values.
@@
For each configuration $C$, define the preference $\textnormal{{pref}}_{p}(C)$ of $p$ in $C$ to be the output value of $p$ in the execution $𝐶𝛼$ where only $p$ takes steps before it decides. We will call such an extension $α$ $p$’s solo-terminating extension and write $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cα)$ to indicate that $p$ decides this value in configuration $𝐶𝛼$. We know that a unique such extension $α$ exists because $p$ is deterministic (making $α$ unique) and the system is wait-free (so that no fairness requirement prevents $p$ from running alone).
@@
Let $α$ be $p$’s solo-terminating extension of $𝐶𝜋$. Then $πα$ is a $p$’s solo-terminating extension of $C$, and $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cπα)=\textnormal{{pref}}_{p}(Cπ)$. ∎
@@
For any wait-free shared-memory protocol satisfying termination and validity, there exists an initial solo-bivalent configuration for any $n≥2$.
@@
$\textnormal{{pref}}_{p}(Cπ_{q})$ $\textnormal{{pref}}_{p}(C)$
@@
$\textnormal{{pref}}_{q}(Cπ_{p})$ $\textnormal{{pref}}_{q}(C).$
@@
5 Hidden threshold conciliator
The ratio between the $O(t)$ time complexity of each process in Theorem 4.1 and the $O(st)$ time complexity of the adversary suggests that a useful strategy for the processes faced with a computationally-limited adversary is to exhaust the adversary’s resources by forcing it to simulate a large number of expensive potential executions.
In this section, we describe how to implement a derandomized conciliator for $n$ processes that reduces the problem of achieving agreement on the existence of a cryptographic implementations of a threshold function where the threshold can only be identified by probing specific function values, which we construct in Section 6. The cost to the adversary to find a bad execution of this conciliator is similar to the upper bound shown in Theorem 4.1.
@@ 16       $i\textleftarrow\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
Algorithm 5.1 Derandomized threshold conciliator
The pseudocode for this conciliator is given in Algorithm 5.1. The main procedure conciliator takes as arguments a consensus input value $\textnormal{{input}}∈\left\{{0,1}\right\}$, a representation of a hidden threshold function $f$, and a security parameter $s$ that scales the number of shared-memory steps taken by the process.
The conciliator proceeds in two phases.
@@
The approximate agreement protocol is an adaptation of Moran’s one-dimensional midpoint protocol [16]. In Moran’s original protocol, each process takes a sequence of snapshots of current values and adopts the midpoint of the values it sees until all fit within a particular range. In our protocol, we organize this sequence into a layered execution of $s$ rounds, where in each round, a process adopts the average value it sees, scaled by a factor of $2$ per round to track the integer numerators instead of the actual fractional values. By taking advantage of the inputs starting at 00 and $1$, we can show that at most two values, differing by at most $1$, appear at the end of each round; this allows us to replace the snapshot with a pair of registers, one of which holds the even value and one the odd. Formally:
@@
Let $S_{r}$ be the set of all $i$ values that are held by any process after $r$ iterations of the loop in procedure approximateAgreement. Then $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$ for some $v_{r}$.
@@
By induction on $r$. For $r=0$, we have $S_{0}⊆\left\{{0,1}\right\}$.
For $r+1$, from the induction hypothesis, there is a value $v_{r}$ such that $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$. So, each process writes either $v_{r}$ or $v_{r+1}$ to one of the registers $A[r][-]$ before reading the other register; since $v_{r}$ and $v_{r}+1$ have different values mod $2$, they will not overwrite each other.
It follows that whichever value $v_{r}$ or $v_{r+1}$ is written first will be visible to all processes. If this is $v_{r}$, then every process sees either $v_{r}$ alone and chooses a new value $2⋅v_{r}$), or sees $v_{r}$ and $v_{r+1}$ and chooses $v_{r}+(v_{r}+1)=2⋅v_{r}+1$. In this case, we get $S_{r+1}⊆\left\{{v_{r+1},v_{r+1}+1}\right\}$ where $v_{r+1}=2⋅v_{r}$. The case where $v_{r}+1$ is written first is similar, except now $v_{r+1}=2⋅v_{r}+1$ instead. ∎
In particular, the set of output values is given by $S_{s}⊆\left\{{v_{s},v_{s}+1}\right\}$ for some $v_{s}$.

### v2

# Obfuscated Consensus

@@ Abstract
The classic Fischer, Lynch, and Paterson impossibility proof FischerLP1985 ; LouiA1987 demonstrates that any deterministic protocol for consensus in either a message-passing or shared-memory system must violate at least one of termination, validity, or agreement in some execution. But it does not provide an efficient procedure to find such a bad execution.
We show that for wait-free shared memory consensus, given a protocol in which each process performs at most $s$ steps computed with total time complexity at most $t$, there exists an adversary algorithm that takes the process’s programs as input and computes within $O(st)$ time a schedule that violates agreement. We argue that this bound is tight assuming the random oracle hypothesis: there exists a deterministic obfuscated consensus protocol that forces the adversary to spend $Ω(st)$ time to find a bad execution despite having full access to all information available to the protocol.
This bound is based on a general algorithm that reduces the constructing an obfuscated consensus protocol to constructing an obfuscated threshold function that provably costs $Ω(t)$ time to evaluate on a single input, where $t$ is a tunable parameter, and for which an adversary with access to the threshold function implementation cannot extract the threshold any faster than by doing binary search. We give a particular implementation of such an obfuscated threshold function that is not very efficient but that is provably secure assuming the random oracle hypothesis. Since our obfuscated consensus protocol does not depend on the specific details of this construction, it may be possible to replace it with one that is more efficient or requires weaker cryptographic assumptions, a task we leave for future work.
@@ 1 Introduction
Asynchronous consensus is a building block in structuring replicated state machines and efficient proof of authority/stake Blockchains, e.g., DBLP:journals/tdsc/DolevGNW24 .
@@
Unfortunately, the celebrated result of Fischer, Lynch, and Paterson FischerLP1985 proves that there is no deterministic algorithm that can yield an asynchronous consensus. Subsequent consensus protocols have escaped this bound by adjusting the requirements of the problem. Paxos DBLP:journals/tocs/Lamport98 succeeded in ensuring the consistency (all participants decide on the same value) and validity (the decided value appears in at least one of the inputs) of the distributed decision but sacrifices the liveness as an adversarial scheduler can keep the system undecided forever. Alternatively, randomization (for example, DBLP:conf/podc/Ben-Or83 ) allows consensus to be solved at the cost of replacing a deterministic termination guarantee with a probabilistic one, where the protocol finishes with probability $1$ in the limit. In these protocols, the adversary scheduler that could otherwise prevent agreement following the procedure given by the FLP proof is unable to predict the future and thus falls off the bad path.
We investigate an alternative approach, where validity and termination are guaranteed in all executions, and agreement holds for all but a small number of executions that are computational expensive to compute, even for an adversary given access to the entire code and knowledge of the processes in the protocol. This idea is inspired by the Fiat-Shamir heuristic FS for extracting a zero-knowledge challenge from a hash of public values assuming a random oracle. In our approach, the scheduling choices of the adversary in an asynchronous shared-memory system, as observed by the processes, become the input to an obfuscated threshold function. Each process computes its output by computing a single value of this function, while the adversary can only prevent agreement through the much more expensive process of finding the threshold input at which the value of the function changes from 0 to $1$. This gap can in principle be exploited to obtain a fully-determinsitic wait-free asynchronous protocol that solves consensus in most executions assuming the adversary has limited computational power.
This approach has some limitations. We show that, given any protocol in which each process performs $s$ steps and computes for $t$ time, there is an adversary program that outputs an agreement-violating schedule in $O(st)$ time in the worst case.
Conversely, we provide a reduction from our obfuscated consensus problem to constructing an obfuscated threshold function, and give a particular implementation of such a function as an obfuscated truth table. Here each bit of the truth table is hidden using a post-quantum time lock puzzle (see, e.g., rivest1996time ; DBLP:conf/asiacrypt/AfsharCHLM23 and the references therein) for efficiently structuring obfuscated programs (see, e.g., DBLP:journals/cacm/JainLS24 and the references therein). This allows us to tune the base cost of computing the threshold function for both the processes and the adversary, and to prevent the adversary from finding the hidden threshold any faster than by binary search over a range exponential in the step complexity of the consensus protocol.
While our particular implementation of an obfuscated threshold function involves a relatively expensive setup phase, it is possible to imagine an implementation using more sophisticated cryptographic techniques that would avoid this overhead. We discuss this possibility further in Section˜7.
1.1 Overview
The system settings are described in the next section. The problem of obfuscated consensus is described in Section 3: this replaces the usual agreement condition for consensus that requires all processes output the same value in each execution with a new obfuscated agreement condition that only requires that it is computationally expensive for the adversary scheduler to find a schedule that violates agreement. Section 5 gives a reduction from obfuscated consensus to a purely cryptographic problem, that of obfuscating a threshold function. This reduction has parameters allowing both the computation cost incurred by each process and the ratio of the adversary’s cost to each process’s cost to be adjusted independently. An implementation of an obfuscated threshold function, along with a proof of security, is given in Section 6. Finally, we discuss some of the limitations of our current implementation and possibilities for future in Section 7.
2 Model
We use a standard asynchronous shared-memory model, with some small tweaks to include explicitly the programs used by the processes and adversary to compute their choices.
@@ 2 Model
The system consists of $n$ asynchronous processes $p_{1},...,p_{n}$ that communicate by reading and writing to shared objects, with timing controlled by an adversary scheduler.
The behavior of each process is controlled by a random-access machine program $P$ that implements a function $Q×V\textrightarrow Q×A$, where $Q$ is the set of states of the process, $V$ is the set of possible return values of operations (including a special null value used to indicate the first step by the process), and $A$ is the next operation to execute.
The adversary scheduler is a program in the same model that takes as input a step bound $s$, a sequence of programs $\left\langle{P_{1},\dots,P_{n}}\right\rangle$ and initial states $q_{1},\dots,q_{n}$, and outputs a a sequence of process ids, where each id occurs at most $s$ times, specifying the order in which the processes take steps. We will refer to this sequence of process ids as a schedule.
Communication is via shared memory: A collection of atomic read/write registers $r_{1},\dots,r_{m}$, where applying a read operation to a register returns its current value and applying a write operations sets the value and returns nothing.
A configuration of the system consists of the states of the processes and registers, together with at most one pending operation for each process.
Every schedule yields an execution, an alternating sequence of configurations and operations $C_{0}α_{1},C_{1}α_{2}C_{2}\dots$ in which each $C_{i+1}$ is the result of applying operation $α_{i+1}$ to configuration $C_{i}$. Because the processes are deterministic, the adversary can simulate the execution corresponding to any particular schedule in time proportional to the total time used by the processes to choose their steps in that execution.
@@
We assume that both the processes and the adversary have access to a random oracle $H$, which is an idealized version of the cryptographic hash function. The random oracle acts as a black box that can be queried with input $x$ to return output $y$. The box is consistent; that is, any input $x$ will always output the same $y$. The box is uniformly random, which means the probability of getting output $y$ for an input $x$ that was not already queried is uniform. Though everyone has access to this black box and can query it, its internal workings are unknown (See Chapter 5 of KatzLindell ).
3 Obfuscated consensus
@@
The Fischer-Lynch-Paterson (FLP) impossibility result FischerLP1985 , as extended to shared memory by Loui and Abu-Amara LouiA1987 , shows that in an asynchronous system with n2n≥2 processes in which at least one process can fail by crashing, there is always an adversary strategy that finds an execution in which at least one of the three conditions is violated. We will be interested in constructing protocols where finding this strategy is computationally difficult.
Define an obfuscated consensus prototocol to be a protocol that satisfies termination and validity always, but replaces agreement with a condition that we can describe informally as
Obfuscated agreement
The cost to the adversary to compute a schedule that violates agreement substantially exceeds the cost to the processes to execute the protocol.
More formally, the adversary must solve an obfuscated consensus problem: Given process programs $P_{1},\dots,P_{n}$ and a step bound $s$, find a schedule in which each process takes $s$ steps that violates one of termination, validity, or agreement. A successful obfuscated consensus protocol is one for which solving this problem is expensive, for some reasonable definition of expensive. We will start by bounding this cost from above.
4 Bounding the cost of finding a bad execution
@@
The core idea of the FLP impossibility proof FischerLP1985 is the notion of bivalence. A configuration $C$ is bivalent if there is a path from it to a configuration where a process decides on the value $1$, and there is also a path from it to another configuration where a process decides on the value 0. Similarly, a configuration $C$ is univalent if all the paths from $C$ lead to configurations with an identical decision value. Because of the agreement property, any configuration in which some process has decided is necessarily univalent.
@@
The FLP proof, in both its original form FischerLP1985 , and its extension to shared memory by Loui and Abu-Amara LouiA1987 , gives an explicit procedure for constructing an infinite non-terminating execution for any protocol that satisfies agreement and validity, by starting in a bivalent configuration and always choosing bivalent successor configurations. However, the computational cost of this construction is not considered. We would like to find an adversary strategy that allows it to prevent a protocol from terminating without having to explore the space of all executions.
@@
The time complexity of computing a non-terminating, and thus infinite, execution is necessarily infinite. We will instead look at the more tractable problem of computing an extension for some fixed number of steps $s$, which is the justification for defining obfuscated consensus as satisfying termination and validity always but possibly failing agreement. As described in Section 2, we replace an abstract adversary function with a concrete adversary program that takes as input the starting configuration of the system and the programs used by the processes to choose their next operations and, ultimately, their outputs.
@@
Formally, let us define the following problem. We assume that each program $P_{i}$ has two special return actions $r_{0}$ and $r_{1}$, that it must choose between after exactly $s$ steps; these provide the output of each process in the consensus protocol. The adversary, a program that takes as input the programs $P_{1},\dots,P_{n}$ and the step bound $s$ and constructs a schedule in which each process takes exactly $s$ steps, wins if any of the processes’ outputs disagree.
@@
Given inputs $P_{1},\dots,P_{n}$ representing processes in a shared-memory system where, in all executions, (a) each process $p_{i}$ outputs a decision value in $s$ steps, and (b) each process $p_{i}$ uses a total of at most $t$ time to compute its transitions, there is an adversary program that computes in $O(st)$ time a schedule that causes some pair of processes to output different decision values.
@@
For each configuration $C$, define the preference $\textnormal{{pref}}_{p}(C)$ of $p$ in $C$ to be the output value of $p$ in the execution Cα where only $p$ takes steps before it decides. We will call such an extension $α$ $p$’s solo-terminating extension and write $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cα)$ to indicate that $p$ decides this value in configuration Cα. We know that a unique such extension $α$ exists because $p$ is deterministic (making $α$ unique) and the system is wait-free (so that no fairness requirement prevents $p$ from running alone).
@@
Let $α$ be $p$’s solo-terminating extension of Cπ. Then $πα$ is a $p$’s solo-terminating extension of $C$, and $\textnormal{{pref}}_{p}(C)=\textnormal{{decision}}_{p}(Cπα)=\textnormal{{pref}}_{p}(Cπ)$. ∎
@@
For any wait-free shared-memory protocol satisfying termination and validity, there exists an initial solo-bivalent configuration for any n2n≥2.
@@
$\textnormal{{pref}}_{p}(Cπ_{q})$ $≠\textnormal{{pref}}_{p}(C)$
@@
$\textnormal{{pref}}_{q}(Cπ_{p})$ $≠\textnormal{{pref}}_{q}(C).$
@@
5 Obfuscated consensus from obfuscated threshold
The ratio between the $O(t)$ time complexity of each process in Theorem 4.1 and the $O(st)$ time complexity of the adversary suggests that a useful strategy for the processes faced with a computationally-limited adversary is to exhaust the adversary’s resources by forcing it to simulate a large number of expensive potential executions. We can do this by making the outcome of each process’s execution depend on evaluating an expensive function of the view that process has in the execution.
An obfuscated threshold function $f:\left\{{0,\dots,\ell}\right\}\textrightarrow\left\{{0,1}\right\}$ is a function, represented as a random-access machine program, with $f(0)=0$ and $f(\ell)=1$. The obfuscated threshold problem is to find, given a representation of such a function, an input $v$ such that f(v)f(v+1)f(v)≠f(v+1). We will show that finding a bad execution of a particular length of any obfuscated consensus protocol is equivalent to solving this problem.
Binary search will find a $v$ with f(v)f(v+1)f(v)≠f(v+1) in $O(\log\ell)$ evaluations of $f$, which puts an upper bound on the cost of solving the obfuscated threshold function problem. But it may be that a sufficiently clever algorithm can extract $v$ from $f$ without evaluating $f$ explicitly. We will discuss how to build an $f$ that resists such attacks in Section 6.3.
Note that the end conditions $f(0)=0$ and $f(1)=1$ show that at least one transitional value $v$ always exists, and, for some choices of $f$, many such transitional values exist. This makes the name “obfuscated threshold function” a bit misleading, since $f$ is not necessarily a threshold function. But choosing a threshold function in particular reduces the likelihood of guessing a bad $v$ at random, so we will try to produce threshold functions if we can.
We will show a reduction between these problems by constructing an obfuscated consensus protocol that uses an obfuscated threshold function to compute its output values. In Algorithm 5.1, we give a protocol that assumes that every process (and thus also the adversary) has access to an obfuscated threshold function $f$, and show that the adversary computes a bad execution for this protocol if and only if it can find a $v$ such that f(v)f(v+1)f(v)≠f(v+1).
@@ 16 $i\textleftarrow\textnormal{{approximateAgreement}}(\textnormal{{input}},s)$
Algorithm 5.1 Obfuscated consensus based on an obfuscated threshold
Pseudocode for a protocol that does this is given in Algorithm 5.1. The main procedure obfuscatedConsensus takes as arguments a consensus input value $\textnormal{{input}}∈\left\{{0,1}\right\}$, a representation of a function $f$ that converts integer values to Boolean decisions, and a security parameter $s$ that scales the number of shared-memory steps taken by the process.
The protocol proceeds in two phases.
@@
The approximate agreement protocol is an adaptation of Moran’s one-dimensional midpoint protocol Moran1995 . In Moran’s original protocol, each process takes a sequence of snapshots of current values and adopts the midpoint of the values it sees until all fit within a particular range. In our protocol, we organize this sequence into a layered execution of $s$ rounds, where in each round, a process adopts the average value it sees, scaled by a factor of $2$ per round to track the integer numerators instead of the actual fractional values. By taking advantage of the inputs starting at 0 and $1$, we can show that at most two values, differing by at most $1$, appear at the end of each round; this allows us to replace the snapshot with a pair of registers, one of which holds the even value and one the odd. Formally:
@@
Let $S_{r}$ be the set of all $i$ values that are held by any process after $r$ iterations of the loop in procedure approximateAgreement. Then $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$ for some $0≤v_{r}≤2^{r}-1$.
@@
By induction on $r$. For $r=0$, we have $v_{0}=0$ and $S_{0}⊆\left\{{0,1}\right\}$.
For $r+1$, from the induction hypothesis, there is a value $v_{r}$ such that $S_{r}⊆\left\{{v_{r},v_{r}+1}\right\}$. So, each process writes either $v_{r}$ or $v_{r+1}$ to one of the registers A[r][]A[r][-] before reading the other register; since $v_{r}$ and $v_{r}+1$ have different values mod $2$, they will not overwrite each other.
It follows that whichever value $v_{r}$ or $v_{r+1}$ is written first will be visible to all processes. If this is $v_{r}$, then every process either sees $v_{r}$ alone and chooses a new value $2⋅v_{r}$, or sees $v_{r}$ and $v_{r+1}$ and chooses $v_{r}+(v_{r}+1)=2⋅v_{r}+1$. In this case, we get $S_{r+1}⊆\left\{{v_{r+1},v_{r+1}+1}\right\}$ where $v_{r+1}=2⋅v_{r}$. The case where $v_{r}+1$ is written first is similar, except now the processes all choose either $v_{r+1}=2⋅v_{r}+1$ or $v_{r+1}+1=2⋅(v_{r}+1)$. ∎
In particular, the set of output values is given by $S_{s}⊆\left\{{v_{s},v_{s}+1}\right\}$ where $0≤v_{s}≤2^{s}-1$.